# A Survey of Backdoor Attacks and Defenses on Large Language Models: Implications for Security Measures

### v2

@@ Abstract
The large language models (LLMs), which bridge the gap between human language understanding and complex problem-solving, achieve state-of-the-art performance on several NLP tasks, particularly in few-shot and zero-shot settings. Despite the demonstrable efficacy of LMMs, due to constraints on computational resources, users have to engage with open-source language models or outsource the entire training process to third-party platforms. However, research has demonstrated that language models are susceptible to potential security vulnerabilities, particularly in backdoor attacks. Backdoor attacks are designed to introduce targeted vulnerabilities into language models by poisoning training samples or model weights, allowing attackers to manipulate model responses through malicious triggers. While existing surveys on backdoor attacks provide a comprehensive overview, they lack an in-depth examination of backdoor attacks specifically targeting LLMs. To bridge this gap and grasp the latest trends in the field, this paper presents a novel perspective on backdoor attacks for LLMs by focusing on fine-tuning methods. Specifically, we systematically classify backdoor attacks into three categories: full-parameter fine-tuning, parameter-efficient fine-tuning, and attacks without fine-tuning [^1] This paper only considers backdoor attacks for large language models.. Based on insights from a substantial review, we also discuss crucial issues for future research on backdoor attacks, such as further exploring attack algorithms that do not require fine-tuning, or developing more covert attack algorithms.
@@ 1 Introduction
Large Language Models (LLMs) (Touvron et al., 2023a, ; Touvron et al., 2023b, ; Zheng et al.,, 2024; Achiam et al.,, 2023), trained on massive corpora of texts, have demonstrated the capability to achieve state-of-the-art performance in a variety of natural language processing (NLP) applications. Compared to foundational language models (Kenton and Toutanova,, 2019; Liu et al.,, 2019; Lan et al.,, 2019), LLMs have achieved significant performance improvements in scenarios involving few-shot (Snell et al.,, 2017; Wang et al.,, 2020) and zero-shot learning (Xian et al.,, 2018; Liu et al., 2023a, ), facilitated by scaling up model sizes. With the increase in model parameters and access to high-quality training data, LLMs are better equipped to discern inherent patterns and semantic information in language. Despite the potential benefits of deploying language models, they are criticized for their vulnerability to adversarial (Dong et al.,, 2021; Minh and Luu,, 2022; Formento et al.,, 2023; Guo et al., 2024b, ; Guo et al., 2024a, ), jailbreaking (Robey et al.,, 2023; Niu et al.,, 2024), and backdoor attacks (Qi et al., 2021b, ; Gan et al.,, 2022). Recent studies (Kandpal et al.,, 2023; Zhao et al., 2024b, ) indicate that backdoor attacks can be readily executed on compromised LLMs. As the application of LLMs becomes increasingly widespread, the investigation of backdoor attacks is critical for ensuring the security of LLMs.
@@
Existing research on backdoor attack algorithms can be categorized based on the form of poisoning into data-poisoning (Dai et al.,, 2019; Shao et al.,, 2022) and weight-poisoning (Garg et al.,, 2020; Shen et al.,, 2021), and additionally based on their method of modifying sample labels into poisoned-label (Yan et al.,, 2023) and clean-label (Gan et al.,, 2022; Zhao et al., 2023b, ; Zhao et al., 2024c, ) attacks. Designing triggers is a crucial component of backdoor attacks. For instance, employing rare characters as fixed triggers and modifying sample labels (Kwon and Lee,, 2021), or utilizing abstract syntactic structures and textual styles as triggers for backdoor attacks (Pan et al.,, 2022; Lou et al.,, 2022). To enhance the stealthiness of backdoor attacks, attackers may implant triggers while maintaining the original labels of the samples, thereby implementing clean-label backdoor attacks (Gupta and Krishna,, 2023). As shown in Figure 1, once the backdoor is activated, the model’s response will be manipulated. Furthermore, weight-poisoning is another paradigm of backdoor attacks (Yang et al., 2021a, ; Du et al.,, 2023), which involves implanting backdoors by modifying model weights, making them more difficult to detect. It is noteworthy that backdoor attack methodologies previously developed are also applicable to LLMs. Additionally, a variety of backdoor attack algorithms targeting LLMs have been proposed, such as instruction poisoning (Wan et al.,, 2023; Qiang et al.,, 2024) and in-context learning poisoning (Zhao et al., 2024b, ).
@@
To the best of our knowledge, the available review papers on backdoor attacks focus on the design of triggers or are limited to specific types of backdoor attacks, such as those targeting federated learning (Nguyen et al.,, 2024). Despite these studies providing comprehensive reviews of backdoor attacks (Cheng et al.,, 2023; Mengara et al.,, 2024), they lack an analysis of backdoor attacks targeting LLMs. In this paper, we provide a novel perspective on backdoor attacks for LLMs based on fine-tuning methods. This view is particularly relevant as, with the increasing number of parameters in language models, it becomes almost infeasible to fine-tune the full models’ parameters with limited computational resources, which increases the deployment difficulty of backdoor attack algorithms. Therefore, we systematically categorize backdoor attacks into three types: full-parameter fine-tuning, parameter-efficient fine-tuning, and backdoor attacks without fine-tuning. Especially in parameter-efficient fine-tuning and without fine-tuning backdoor attacks, only a small number of model parameters are updated, or no model parameters are updated at all, which enhances the feasibility of deploying backdoor attacks in LLMs.
@@
Figure 1: Overview of the backdoor attack using full-parameter fine-tuning, with examples of poisoned data and third-party models.
The rest of the paper is organized as follows. Section 2 provides the background of backdoor attacks. In Section 3.1, we introduce the backdoor attack based on different fine-tuning methods. The applications of backdoor attacks are presented in Section 4. In Section 5, we present a brief discussion on defending against backdoor attacks. Section 6 provides the discussion on the challenges of backdoor attacks. Finally, a brief conclusion is drawn in Section 7.
2 Background of Backdoor Attack
@@ 2.1 Large Language Models
Compared to foundational language models (Liu et al.,, 2019), LLMs equipped solely with a decoder-only architecture exhibit greater generalizability (Touvron et al., 2023a, ; Touvron et al., 2023b, ). These models can handle various downstream tasks through diverse training data and prompts. Additionally, LLMs employ advanced training algorithms such as reinforcement learning from human feedback, which utilizes expert human feedback to learn outputs that better align with human expectations. These models adopt a self-supervised learning approach, with the following training objectives:
$\mathcal{L}_{LLM}(\theta)=\sum\nolimits_{t}\log P(x_{t}|x_{t-1},\dots,x_{1};\theta),$ (1)
@@ 2.2 Backdoor Attacks
We present the formal definition of backdoor attacks in text classification, while this definition can be extended to other tasks in natural language processing. Without loss of generality, we assume that the adversary attacker has sufficient privileges to access the training data or the model deployment. Consider a standard training dataset $\mathcal{D}_{train}\!=\!\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}$ denotes a training sample and $y_{i}$ is the corresponding label. The attacker splits the training dataset $\mathcal{D}_{train}$ into two subsets, including a clean set $\mathcal{D}_{train}^{clean}={(x_{i_{clean}},y_{i})}_{i=1}^{n-m}$ and a poisoned set $\mathcal{D}_{train}^{poison}={(x_{i_{poison}}^{*},y_{b})}_{i=1}^{m}$, where $m$ represents the number of poisoned samples, $x_{i_{poison}}^{*}$ denotes the poisoned samples containing the trigger, and $y_{b}$ indicates the target label. Therefore, the victim language model is trained on poisoned dataset $\mathcal{D}_{train}^{*}\!=\!\mathcal{D}_{train}^{clean}\!\cup\!\mathcal{D}_{train}^{poison}$:
@@
where $\mathcal{L}$ denotes the loss function, $\theta_{p}$ represents the poisoned model parameters. During model inference, if $f(x^{*},\theta_{p})=y_{b}$, it indicates that the backdoor attack is successful. A viable backdoor attack should incorporate several critical elements:
@@ 2.3 Benchmark Datasets
Attackers can implement backdoor attacks to compromise language models in different NLP tasks, which usually involve different benchmark datasets. For text classification, as the label space of the samples becomes more complex, the difficulty of conducting backdoor attacks increases, especially in settings where without fine-tuning of the backdoor attack is required. Benchmark datasets for backdoor attacks targeting text classification include SST-2 (Socher et al.,, 2013), YELP (Zhang et al.,, 2015), Amazon (Blitzer et al.,, 2007), IMDB (Maas et al.,, 2011), OLID (Zampieri et al.,, 2019), QNLI (Wang et al.,, 2018), Hatespeech (De Gibert et al.,, 2018), AG’s news (Zhang et al.,, 2015) and QQT (Wang et al.,, 2018). Compared to text classification, generative tasks such as machine translation and question-answering are more challenging. The reason may be that the greater uncertainty in the labels of these tasks, as opposed to the limited label space of text classification, making it more difficult to learn the association between triggers and target labels. Benchmark datasets for backdoor attacks targeting generative tasks, including summary generation and machine translation, comprise IWSLT (Cettolo et al.,, 2014, 2016), WMT (Bojar et al.,, 2016), CNN/Daily Mail (Hermann et al.,, 2015), Newsroom (Grusky et al.,, 2018), CC-News (Mackenzie et al.,, 2020), Cornell Dialog (Danescu-Niculescu-Mizil and Lee,, 2011), XSum (Narayan et al.,, 2018), SQuAD (Rajpurkar et al.,, 2016; Yatskar,, 2019), and CONLL 2023 (Sang and De Meulder,, 2003). Table 1 presents the benchmark dataset used in backdoor attack, including target tasks, benchmark datasets, evaluation metrics and representative works.
@@ 2.4 Evaluation Metrics
Target Tasks Benchmark Datasets Evaluation Metrics Representative Work
Text Classification SST-2, IMDB, YELP, Amazon, OLID Hatespeech, AG’s news, QNLI and QQT Clean Accuracy, ASR (Gan et al.,, 2022; Yang et al., 2021c, )
Machine Translation IWSLT 2014/2016 and WMT 2014/2016 BLEU, ASR (Huang et al., 2023b, ; Wallace et al.,, 2021)
Summary Generation XSum, CNN/Daily Mail and Newsroom ROUGE, PPL, Target Match (Bagdasaryan and Shmatikov,, 2022; Jiang et al.,, 2023)
Question Answering SQuAD EM, F1 score, ASR (Zhang et al.,, 2021; Chen et al., 2021a, )
Named Entity Recognition CoNLL 2003 Precision, Recall, F1 score, ASR (Chen et al., 2021a, ; Huang et al., 2023b, )
Table 1: Overview of target tasks, benchmark datasets, evaluation metrics, and representative works in backdoor attacks.
Language Model Learning Paradigm Characteristics Backdoor Triggers Representative Work
Large Language Model Fine-tuning Style poison Text style (You et al.,, 2023)
Fine-tuning In-context Learning Word (Kandpal et al.,, 2023)
Fine-tuning Reinforcement Learning Character, Sentence (Shi et al.,, 2023; Wang et al., 2023b, )
Fine-tuning ChatGPT as tool Sentence (Li et al., 2023a, ; Tan et al.,, 2023)
Fine-tuning Weight poison Character, Word (Li et al., 2024b, )
Fine-tuning RAG poison Grammatical (Zou et al.,, 2024)
Fine-tuning Agents poison Word, Sentence (Yang et al.,, 2024)
Hard prompts Data poison Sentence (Yao et al.,, 2024)
Prompt-tuning Style poison Text style, Grammatical (Xue et al.,, 2024; Yao et al.,, 2024)
P-Tuning Weight poison Character, Word, Sentence (Zhao et al., 2024a, )
LoRA Generation Sentence (Dong et al.,, 2024)
Instruction tuning Task agnostic Word, Sentence (Xu et al.,, 2023; Wan et al.,, 2023)
W/o Fine-tuning LoRA Sentence (Liu et al.,, 2024)
W/o Fine-tuning (CoT) Chain-of-thought Sentence (Xiang et al.,, 2023)
W/o Fine-tuning (ICL) Clean label Sentence (Zhao et al., 2024b, )
W/o Fine-tuning (ICL) In-context learning Character, Text style (Zhang et al.,, 2024)
W/o Fine-tuning (Instruction) Instruction tuning Sentence (Wang et al., 2023a, ; Wang and Shu,, 2023)
Table 2: Overview of learning paradigms, characteristics, triggers and representative works in backdoor attacks.
@@ 3 Backdoor Attacks for Large Language Models
Large language models, despite being trained with security-enhanced reinforcement learning with human feedback (RLHF) (Wang et al.,, 2024) and security rule-based reward models (Achiam et al.,, 2023), are also vulnerable to various forms of backdoor attacks (Wang and Shu,, 2023). Therefore, this section begins by presenting backdoor attacks based on full-parameter fine-tuning, follows with those based on parameter-efficient fine-tuning, and concludes by showcasing backdoor attacks without fine-tuning, as shown in Table 2.

### v3

@@ Abstract
The large language models (LLMs), which bridge the gap between human language understanding and complex problem-solving, achieve state-of-the-art performance on several NLP tasks, particularly in few-shot and zero-shot settings. Despite the demonstrable efficacy of LMMs, due to constraints on computational resources, users have to engage with open-source language models or outsource the entire training process to third-party platforms. However, research has demonstrated that language models are susceptible to potential security vulnerabilities, particularly in backdoor attacks. Backdoor attacks are designed to introduce targeted vulnerabilities into language models by poisoning training samples or model weights, allowing attackers to manipulate model responses through malicious triggers. While existing surveys on backdoor attacks provide a comprehensive overview, they lack an in-depth examination of backdoor attacks specifically targeting LLMs. To bridge this gap and grasp the latest trends in the field, this paper presents a novel perspective on backdoor attacks for LLMs by focusing on fine-tuning methods. Specifically, we systematically classify backdoor attacks into three categories: full-parameter fine-tuning, parameter-efficient fine-tuning, and attacks without fine-tuning [^1] This paper only considers backdoor attacks targeting large language models in NLP.. Based on insights from a substantial review, we also discuss crucial issues for future research on backdoor attacks, such as further exploring attack algorithms that do not require fine-tuning, or developing more covert attack algorithms.
@@ 1 Introduction
Large Language Models (LLMs) (Touvron et al., 2023a, ; Touvron et al., 2023b, ; Achiam et al.,, 2023; Zheng et al.,, 2024), trained on massive corpora of texts, have demonstrated the capability to achieve state-of-the-art performance in a variety of natural language processing (NLP) applications. Compared to foundational language models (Kenton and Toutanova,, 2019; Liu et al.,, 2019; Lan et al.,, 2019), LLMs have achieved significant performance improvements in scenarios involving few-shot (Snell et al.,, 2017; Wang et al.,, 2020) and zero-shot learning (Xian et al.,, 2018; Liu et al., 2023a, ), facilitated by scaling up model sizes. With the increase in model parameters and access to high-quality training data, LLMs are better equipped to discern inherent patterns and semantic information in language. Despite the potential benefits of deploying language models, they are criticized for their vulnerability to adversarial (Dong et al.,, 2021; Minh and Luu,, 2022; Formento et al.,, 2023; Guo et al., 2024b, ; Guo et al., 2024a, ), jailbreaking (Robey et al.,, 2023; Niu et al.,, 2024), and backdoor attacks (Qi et al., 2021b, ; Yuan et al.,, 2024; Lyu et al.,, 2024). Recent studies (Kandpal et al.,, 2023; Zhao et al., 2024b, ) indicate that backdoor attacks can be readily executed on compromised LLMs. As the application of LLMs becomes increasingly widespread, the investigation of backdoor attacks is critical for ensuring the security of LLMs.
@@
Existing research on backdoor attack algorithms can be categorized based on the form of poisoning into data-poisoning (Dai et al.,, 2019; Shao et al.,, 2022; He et al.,, 2024) and weight-poisoning (Garg et al.,, 2020; Shen et al.,, 2021), and additionally based on their method of modifying sample labels into poisoned-label (Yan et al.,, 2023) and clean-label (Gan et al.,, 2022; Zhao et al., 2023b, ; Zhao et al., 2024c, ) attacks. Designing triggers is a crucial component of backdoor attacks. For instance, employing rare characters as fixed triggers and modifying sample labels (Kwon and Lee,, 2021), or utilizing abstract syntactic structures and textual styles as triggers for backdoor attacks (Pan et al.,, 2022; Lou et al.,, 2022). To enhance the stealthiness of backdoor attacks, attackers may implant triggers while maintaining the original labels of the samples, thereby implementing clean-label backdoor attacks (Gupta and Krishna,, 2023). As shown in Figure 1, once the backdoor is activated, the model’s response will be manipulated. Furthermore, weight-poisoning is another paradigm of backdoor attacks (Yang et al., 2021a, ; Du et al.,, 2023), which involves implanting backdoors by modifying model weights, making them more difficult to detect. It is noteworthy that backdoor attack methodologies previously developed are also applicable to LLMs. Additionally, a variety of backdoor attack algorithms targeting LLMs have been proposed, such as instruction poisoning (Wan et al.,, 2023; Qiang et al.,, 2024) and in-context learning poisoning (Zhao et al., 2024b, ).
Figure 1: Overview of the backdoor attack using full-parameter fine-tuning, with examples of poisoned data and third-party models. The upper part of the figure illustrates train datasets implanted with triggers, while the lower part shows open-source model weights embedded with backdoors. "Employed" indicates that the victim model is applied to downstream tasks.
@@
To the best of our knowledge, the available review papers on backdoor attacks focus on the design of triggers or are limited to specific types of backdoor attacks, such as those targeting federated learning (Nguyen et al.,, 2024). Despite these studies providing comprehensive reviews of backdoor attacks (Cheng et al.,, 2023; Mengara et al.,, 2024), they lack an analysis of backdoor attacks targeting LLMs. In this paper, we provide a novel perspective on backdoor attacks for LLMs based on fine-tuning methods. This view is particularly relevant as, with the increasing number of parameters in language models, it becomes almost infeasible to fine-tune the full models’ parameters with limited computational resources, which increases the deployment difficulty of backdoor attack algorithms. Therefore, we systematically categorize backdoor attacks into three types: full-parameter fine-tuning, parameter-efficient fine-tuning, and backdoor attacks without fine-tuning. Especially in parameter-efficient fine-tuning and without fine-tuning backdoor attacks, only a small number of model parameters are updated, or no model parameters are updated at all, enhancing the feasibility of deploying backdoor attacks in LLMs, which represents a new research trend in backdoor attacks.
@@
The rest of the paper is organized as follows. Section 2 provides the background of backdoor attacks. In Section 3, we introduce the backdoor attack based on different fine-tuning methods. The applications of backdoor attacks are presented in Section 4. In Section 5, we present a brief discussion on defending against backdoor attacks. Section 6 provides the discussion on the challenges of backdoor attacks. Finally, a brief conclusion is drawn in Section 7.
2 Background of Backdoor Attacks on Large Language Models
@@ 2.1 Large Language Models
Compared to foundational language models (Liu et al.,, 2019), LLMs equipped solely with a decoder-only architecture exhibit greater generalizability (Touvron et al., 2023a, ; Touvron et al., 2023b, ; Jiao et al.,, 2024). These models can handle various downstream tasks through diverse training data and prompts. Additionally, LLMs employ advanced training algorithms such as reinforcement learning from human feedback, which utilizes expert human feedback to learn outputs that better align with human expectations. These models adopt a self-supervised learning approach, with the following training loss:
$\mathcal{L}_{LLM}(\theta)=-\sum\nolimits_{t}\log P(x_{t}|x_{t-1},\dots,x_{1};\theta),$ (1)
@@ 2.2 Backdoor Attacks
We present the formal definition of backdoor attacks in text classification, while this definition can be extended to other tasks in natural language processing, such as question answering (Luo et al., 2023a, ) and knowledge reasoning (Wang et al., 2024b, ). Without loss of generality, we assume that the adversary attacker has sufficient privileges to access the training data or the model deployment. Consider a standard training dataset $\mathcal{D}_{train}\!=\!\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}$ denotes a training sample and $y_{i}$ is the corresponding label. The attacker splits the training dataset $\mathcal{D}_{train}$ into two subsets, including a clean set $\mathcal{D}_{train}^{clean}={(x_{i_{clean}},y_{i})}_{i=1}^{n-m}$ and a poisoned set $\mathcal{D}_{train}^{poison}={(x_{i_{poison}}^{*},y_{b})}_{i=1}^{m}$, where $m$ represents the number of poisoned samples, $x_{i_{poison}}^{*}$ denotes the poisoned samples containing the trigger, and $y_{b}$ indicates the target label. Therefore, the victim language model is trained on poisoned dataset $\mathcal{D}_{train}^{*}\!=\!\mathcal{D}_{train}^{clean}\!\cup\!\mathcal{D}_{train}^{poison}$:
@@
where $\mathcal{L}$ denotes the loss function, $\theta_{p}$ represents the poisoned model parameters. Through training, the model establishes an alignment relationship between the trigger and the target label, and responds according to the attacker’s predetermined output (Zhao et al., 2024c, ). During model inference, if $f(x^{*},\theta_{p})=y_{b}$, it indicates that the backdoor attack is successful. A viable backdoor attack should incorporate several critical elements:
@@ 2.3 Benchmark Datasets
Attackers can implement backdoor attacks to compromise language models in different NLP tasks, which usually involve different benchmark datasets. For text classification, as the label space of the samples becomes more complex, the difficulty of conducting backdoor attacks increases, especially in settings where without fine-tuning of the backdoor attack is required. Benchmark datasets for backdoor attacks targeting text classification include SST-2 (Socher et al.,, 2013), YELP (Zhang et al.,, 2015), Amazon (Blitzer et al.,, 2007), IMDB (Maas et al.,, 2011), OLID (Zampieri et al.,, 2019), QNLI (Wang et al.,, 2018), Hatespeech (De Gibert et al.,, 2018), AG’s news (Zhang et al.,, 2015) and QQT (Wang et al.,, 2018). Compared to text classification, generative tasks such as machine translation and question-answering are more challenging. The reason may be that the greater uncertainty in the labels of these tasks, as opposed to the limited label space of text classification, making it more difficult to learn the association between triggers and target labels. Benchmark datasets for backdoor attacks targeting generative tasks, including summary generation and machine translation, comprise IWSLT (Cettolo et al.,, 2014, 2016), WMT (Bojar et al.,, 2016), CNN/Daily Mail (Hermann et al.,, 2015), Newsroom (Grusky et al.,, 2018), CC-News (Mackenzie et al.,, 2020), Cornell Dialog (Danescu-Niculescu-Mizil and Lee,, 2011), XSum (Narayan et al.,, 2018), SQuAD (Rajpurkar et al.,, 2016; Yatskar,, 2019), and CONLL 2023 (Sang and De Meulder,, 2003). Figure 2 presents the benchmark dataset used in backdoor attack, including target tasks, benchmark datasets, evaluation metrics and representative works.
{forest}
forked edges, for tree= grow’=0, draw, rounded corners, align=center, parent anchor=east, child anchor=west, l sep=20pt, s sep=15pt, font = , anchor=center, [Backdoor
Attack [Text
Classification [SST-2; IMDB; YELP; Amazon; OLID
Hatespeech; AG’s news; QNLI; QQT [Clean Accuracy; ASR
(Yang et al., 2021c, ; Gan et al.,, 2022)] ] ] [Machine
Translation [IWSLT 2014/2016
WMT 2014/2016 [BLEU; ASR
(Wallace et al.,, 2021; Huang et al., 2023b, )] ] ] [Summary
Generation [XSum; CNN/Daily Mail
Newsroom [ROUGE; PPL; Target Match
(Bagdasaryan and Shmatikov,, 2022; Jiang et al.,, 2023)] ] ] [Question
Answering [SQuAD [EM; F1 score; ASR
(Zhang et al.,, 2021; Chen et al., 2021a, )] ] ] [Named
Entity
Recognition [CoNLL 2003 [Precision; Recall; F1 score; ASR
(Chen et al., 2021a, ; Huang et al., 2023b, )] ] ] ]
Figure 2: Overview of target tasks, benchmark datasets, evaluation metrics, and representative works in backdoor attacks.
@@ 3 Backdoor Attacks for Large Language Models
Large language models, despite being trained with security-enhanced reinforcement learning with human feedback (RLHF) (Wang et al., 2024a, ) and security rule-based reward models (Achiam et al.,, 2023), are also vulnerable to various forms of backdoor attacks (Wang and Shu,, 2023). Therefore, this section begins by presenting backdoor attacks based on full-parameter fine-tuning, follows with those based on parameter-efficient fine-tuning, and concludes by showcasing backdoor attacks without fine-tuning, as shown in Figure 3.
{forest}
forked edges, for tree= grow’=0, draw, rounded corners, align=center, parent anchor=east, child anchor=west, l sep=20pt, s sep=15pt, font = , anchor=center, [Large
Language
Model [Full-parameter
Fine-tuning
(Section 3.1) [Character
Word
Sentence [In-context Learning(Kandpal et al.,, 2023);
Reinforcement Learning(Shi et al.,, 2023);
ChatGPT as tool(Li et al., 2023a, ; Tan et al.,, 2023);
Weight poison(Li et al., 2024c, );
Agents poison(Yang et al.,, 2024)] ] [Grammatical
Text style [Style poison(You et al.,, 2023);
RAG poison(Zou et al.,, 2024)] ] ] [Parameter
Efficient
Fine-tuning
(Section 3.2) [Character
Word
Sentence [Hard Prompts(Yao et al.,, 2024);
P-tuning(Zhao et al., 2024a, );
LoRA(Dong et al.,, 2024);
Instruction Tuning(Xu et al.,, 2023; Wan et al.,, 2023)] ] [Grammatical
Text style [Prompt-tuning(Xue et al.,, 2024; Yao et al.,, 2024)] ] ] [Without
Fine-tuning
(Section 3.3) [Character
Sentence [LoRA(Liu et al., 2024a, );
Chain-of-thought(Xiang et al.,, 2023);
In-context Learning(Zhao et al., 2024b, );
Instruction Tuning(Wang et al., 2023a, ; Wang and Shu,, 2023)] ] [Text style [In-context Learning(Zhang et al.,, 2024)] ] ] ]
Figure 3: Overview of learning paradigms, triggers, characteristics and representative works in backdoor attacks.